# RPBG: Towards Robust Neural Point-based Graphics in the Wild

### v1

@@ 1 Introduction
For thorough qualitative and quantitative evaluation, we collect 4 typical challenging scene types as the benchmark for robust NVS, i.e., 360° unbounded scenes (with free trajectories), inside-out scenes, large-scale scenes (at the scale of a block or campus), and sparse-view scenes. As a result, RPBG exhibits great robustness with perceptually satisfactory synthesis across the aforementioned typical scenes types as showcased in Fig. 1, where the high-quality re-renderings are obtained with an exactly identical set of hyper-parameters without any manual configuration, relieving the practitioners from the exhaustive per-scene search of hyper-parameters. Its stable superiority over state-of-the-art NVS methods, we believe, is of great significance especially for real applications.
@@ 2 Related Work
In contrast to conventional physically based rendering (PBR) techniques, neural rendering [46] learns to render high-quality images in a data-driven manner. Particularly, we focus on the methods that conduct neural re-rendering with point clouds. Bui et al. [5] propose to enhance the coarse point-based rendering by a GAN for image super-resolution. NRW [29] and InvSFM [33] attempt to re-render the reconstructed point cloud with respective auxiliary buffers, e.g., latent appearance vectors, semantic masks and SIFT descriptors [27]. NPBG [1, 34] adopts a U-Net-like [37] CNN to render neural point textures as RGB images, and exhibits better flexibility over mesh-based proxies [47]. The practical applications in the context of autonomous driving, e.g., scene editing and stitching, and large-scale training, are further explored in [25]. Recently, ADOP [38] and Gaussian Splatting [22] have demonstrated notable accomplishments. However, the differentiable rendering/splatting scheme employed requires an extensive amount of memory to maintain the computational graph and gradients during training. This limitation hinders their application on large-scale scenes.
@@ 3.1 Preliminaries
NPBG [1] and NPBG++ [34] apply a non-differentiable hard point $z$-buffering operation as an approximated back-face culling when rasterizing points as 2D fragments, where the fragment is updated when and only when the newly projected point has a smaller $z$-depth than the current one. After a traversal of all the points, the fragment keeps the record of the indices of rasterized points. Then tensor scattering is performed to sample the neural texture at the corresponding positions. Compared to differentiable rendering/splatting alternatives [38, 22], we consider the most significant advantage is its memory efficiency and thus better scalability, since the computation graph required to keep is much smaller for NPBG. We also want to keep this scalability while avoiding interfering the elegant rasterization paradigm.
@@
The renderer $\mathcal{R}_{\Theta}:\mathbb{R}^{H\times W\times C}\rightarrow\mathbb{R}^{H\times W\times 3}$ of NPBG [1] follows a vanilla U-Net-like architecture [37]. Since the target point clouds are supposed to be is well-constructed points, the vanilla U-Net is expressive enough to complete the assigned task of mapping higher dimensional features ($C$ set as 8) from the neural texture $\mathbf{T}$ to RGB values.
@@ 4.2 Point Triangulation
Note that this strategy is optional, and we only apply such strategy to scenes with extremely poor triangulation. Please refer to the Supplementary Material for more information.
@@ 5.2 Implementation Details
We randomly crop images to square patches of $256\times 256$, with a batch size of 8. The learning rate for the neural point textures is $10^{-1}$, and $10^{-4}$ for the rendering networks, which will decay by a factor of 0.5 if 5 consequent epochs witness no drop in the loss function. A dimension of 8 is applied for any neural texture regardless of the scene scale or complexity. For the weights of loss functions, we globally set $\lambda_{\mathrm{huber}}=10^{3}$, $\lambda_{\mathrm{vgg}}=1$, and $\lambda_{\mathrm{fft}}=1$. The training is performed on one NVIDIA GeForce RTX 3090, with a GPU memory consumption of up to 23 GB. It takes around 8 to 30 GPU hours for training. We would like to emphasize that, RPBG is equipped with the exactly identical set of hyper-parameters or tuning strategies for ALL the scenes covered by the experiments.
@@ 5.4 Ablation Study
We study different strategies of environment modeling for NVS in-the-wild, namely leaving the blank pixels with zeros, filling the blanks with a default learnable feature (as is done in RPBG), and wrapping the points with a sphere of $10^{6}$ points (equivalent to an environment map). The quantitative results are in Tab. 3, where the strategy incorporated by RPBG boosts the performance by a large margin, with almost no additional overhead.
@@
Table 3: Ablation on environment modeling on two unbounded scenes, namely Caterpillar of T&T dataset [23] and sky of Free dataset [52]. #Points represent the equivalent overheads when applying different strategies.
@@
Zeros 7,952,015 19.89 0.456 0.403 17,489,399 21.77 0.598 0.315
Learnable 7,952,016 21.78 0.687 0.289 17,489,400 24.81 0.866 0.199
Sphere 8,952,015 21.68 0.686 0.283 18,489,399 25.15 0.869 0.179
@@ 6 Discussion
We would like to shed some light on the effectiveness of RPBG. The first factor should be the point representation. The triangulated 3D points have already contained all the verified co-visibility that is required for NVS. Compared to RF-based methods, which often struggle at local minima, such triangulation-as-parameterization paradigm greatly lower the complexity of searching, and is unified across different scene types. Besides, the patch-wise rendering scheme mentioned in Sec. 4.4 also improves the perceptual quality of re-rendering while RF-based methods apply pixel/ray-wise volume rendering, where the correlation between pixels are not modeled explicitly. For more discussions on the insights, please refer to the Supplementary Material.
@@
An obvious problem with point-based re-rendering (RPBG and NPBG [1]) is that compared to lightweight RF-based variants [7, 31], they take more space to store the CNN parameters and the neural texture. It also makes the scaling-up easier, that the number of parameters grows linearly with the point cloud scale, while the RF-based methods require explicit partition protocol [44, 48]. Besides, since the non-local neural renderer we employ, each point is encoded with visual context of a larger range, RPBG’s editability is decreased. Please refer to the Supplementary Material for analysis.
@@ 7 Conclusion
In this paper, we present RPBG as a robust and practical alternative to NPBG, a baseline of point-based NVS methods, performing neural re-rendering on triangulated points. We analyze the key problems in NPBG, when attempting to generalize to more generic scenes other than the well-captured scans, and reform the pipeline to reveal the real potential of point-based graphics. Respectively motivated and inspired by RF-based methods and low-level image restoration methods, we reform the pipeline according to our analysis. By extensive experiments on diverse datasets, RPBG achieves stably superior results over state-of-the-art RF-based and point-based NVS methods, especially on the metrics with more attention paid to the nuances of human visual perception, with the exactly identical set of hyper-parameters across ALL scenes, indicating its robustness and generalizability.
@@ Appendix 0.A Further Discussions
As the framed area in Fig. 9(a) shows, the rendering noise is mainly caused by under-representation of RF, which is further due to the sparsity of input views (lack of ray intersections). RF-based methods aim to represent the target scene loyally, where each inquiry is supposed to be a frank reflection of local optical properties. In this way, RF-based methods render an image in pixels without considering the context information, establishing better pixel-to-pixel correspondence (higher PSNR).
@@
Figure 9: Comparison on Museum of T&T dataset [23] to showcase the different sources of noise by the fundamental differences between methods. Zoom in for best view.
@@ Appendix 0.B Dense Triangulation
Based on the sparse triangulation, we follow the view selection strategy in [58], and choose 4 neighboring images with the best co-visibility for each image. Then we estimate a depth map for each image, by aid of the 4 neighboring images, with AA-RMVSNet [54]. The per-view depth maps are filtered and fused to obtain the final 3D point cloud. We select AA-RMVSNet for its high memory-efficiency that allows a large batch size and RPBG is supposed to work fine with other off-the-self MVS methods.
@@ Appendix 0.C Point Cloud
The detailed augmentation steps are as follows.
  1. 1.
    Sample one existing point $X=(x,y,z)$ randomly from the point cloud, and form a 3D Gaussian distribution $G$ with its mean value as $X$.
  2. 2.
    Sample one new point $X^{{}^{\prime}}$ from $G$ and add this point to the point cloud.
  3. 3.
    Repeat (1)-(2) for a given number of times.
  4. 4.
    Train RPBG with the new point cloud.
  5. 5.
    After training, we validate the newly added points by thresholding the pseudo density defined as $\sigma_{i}=\sum|\mathbf{T}_{i}|$. Points with lower $\sigma$ are pruned.
On the scene of Church, we study the impact of such strategy applied on sparsely triangulated points multiple times. As is shown in Tab. 4, the first round of sampling and pruning brings the largest performance gain, and a larger gain is observed when applying to the SfM-initialized triangulation, which is much sparser compared to the MVS-initialized one.
@@
Since we have demonstrated in the main paper that RPBG is able to perform re-rendering even with a randomly initialized point cloud taken as input. Empirically, we find that by applying the spatial pruning strategy to the random point cloud (by thresholding point-wise $\sigma>180$), the point cloud shrinks to a shape similar to the actual geometry, as is illustrated in Fig. 10(b). It suggests that when neurally re-rendering, the network is able to implicitly verify the occupancy of each rasterized point and if a point is observed with poor multi-view consistency, it is more likely to be considered as an invalid point. The attempt of pruning random points is considered as an extreme case explaining how the point cloud augmentation strategy of RPBG manages to alleviate the problem of patchy or erroneous triangulation.
@@
The RF-based methods are sensitive to dynamic objects and require non-trivial data pre-processing, e.g., masking by either manual labeling or semantic segmentation, to prevent ambiguity in optimization. As for RPBG, since dynamic objects do not satisfy the static scene assumption, they are typically not reconstructed in SfM or MVS. By experiments, we discover that RPBG is robust to such dynamic objects and able to automatically such objects in the training views when re-rendering (Fig. 11), which suggests that multi-view consistency is implicitly enforced during training and the renderer tends to restore the most consensual re-rendering.
@@
As RPBG is a point-based pipeline, where an explicit 3D geometry is adopted for re-rendering, it enables scene editing and inpainting to a certain extent. We give two examples in Fig. 12. By removing the points, together with the point-bounded features, RPBG manages to re-render the edited scene, yet with some artifacts observed. It is because in RPBG, we enhance the context exchange among rasterized points by DAC, where each point does not solely represent its local optical property.
@@ Appendix 0.D More Quantitative Results
ScanNet++ Dataset
We also evaluate RPBG on the public benchmark of ScanNet++ [59] (Novel View Synthesis on DSLR Images). The results are shown in Tab. 7. Note that the scores are all retrieved from the leaderboard. RPBG outperforms all the baselines listed by the benchmark, with a particular good perceptual quality (LPIPS).

### v2

@@ 1 Introduction
For thorough qualitative and quantitative evaluation, we collect 4 typical challenging scene types as the benchmark for robust NVS, i.e., 360° unbounded scenes (with free trajectories), inside-out scenes, large-scale scenes (at the scale of a block or campus), and sparse-view scenes. As a result, RPBG exhibits great robustness with perceptually satisfactory synthesis across the aforementioned typical scenes types as showcased in Fig. 1, where the high-quality re-renderings are obtained with an exactly identical parameterization strategy without any manual configuration, relieving the practitioners from the exhaustive per-scene search of hyper-parameters. Its stable superiority over state-of-the-art NVS methods, we believe, is of great significance especially for real applications.
@@ 2 Related Work
In contrast to conventional physically based rendering (PBR) techniques, neural rendering [48] learns to render high-quality images in a data-driven manner. Particularly, we focus on the methods that conduct neural re-rendering with point clouds. Bui et al. [5] propose to enhance the coarse point-based rendering by a GAN for image super-resolution. NRW [30] and InvSFM [34] attempt to re-render the reconstructed point cloud with respective auxiliary buffers, e.g., latent appearance vectors, semantic masks and SIFT descriptors [28]. NPBG [1, 35] adopts a U-Net-like [38] CNN to render neural point textures as RGB images, and exhibits better flexibility over mesh-based proxies [49]. The practical applications in the context of autonomous driving, e.g., scene editing and stitching, and large-scale training, are further explored in READ [26]. Recently, ADOP [39] and Gaussian Splatting [23] have demonstrated notable accomplishments. However, the differentiable rendering/splatting scheme employed requires an extensive amount of memory to maintain the computational graph and gradients during training. This limitation hinders their application on large-scale scenes.
@@ 3.1 Preliminaries
NPBG [1] and NPBG++ [35] apply a non-differentiable hard point $z$-buffering operation as an approximated back-face culling when rasterizing points as 2D fragments, where the fragment is updated when and only when the newly projected point has a smaller $z$-depth than the current one. After a traversal of all the points, the fragment keeps the record of the indices of rasterized points. Then tensor scattering is performed to index the neural texture at the corresponding positions. Compared to differentiable rendering/splatting alternatives [39, 23], we consider the most significant advantage is its memory efficiency and thus better scalability, since the computation graph required to keep is much smaller for NPBG. We also want to keep this scalability without interfering the elegant rasterization paradigm.
@@
The renderer $\mathcal{R}_{\Theta}:\mathbb{R}^{H\times W\times C}\rightarrow\mathbb{R}^{H\times W\times 3}$ of NPBG [1] follows a vanilla U-Net-like architecture [38]. Since the target point clouds for rendering are supposed to be well-constructed, the vanilla U-Net is expressive enough to complete the assigned task of mapping higher dimensional features from the neural texture $\mathbf{T}$ to RGB values. Empirically, we find the expected properties of $\mathcal{R}_{\Theta}$ have much in common with a network for low-level vision tasks, e.g., image restoration.
@@ 4.2 Point Triangulation
Note that this strategy is optional, and we only apply such strategy to scenes with extremely poor triangulation. Please refer to the Supplementary Material for more discussions regarding the self-pruning.
@@ 5.2 Implementation Details
We randomly crop images to square patches of $256\times 256$, with a batch size of 8. The learning rate for the neural point textures is $10^{-1}$, and $10^{-4}$ for the rendering networks, which will decay by a factor of 0.5 if 5 consequent epochs witness no drop in the loss function. A dimension of 8 is applied for any neural texture regardless of the scene scale or complexity. For the weights of loss functions, we globally set $\lambda_{\mathrm{huber}}=10^{3}$, $\lambda_{\mathrm{vgg}}=1$, and $\lambda_{\mathrm{fft}}=1$. The training is performed on one NVIDIA GeForce RTX 3090, with a GPU memory consumption of up to 23 GB. It takes around 8 to 30 GPU hours for training, depending on the data scale. We would like to emphasize that, RPBG does not require case-by-case scene parameterization or grid search of training hyper-parameters for all the scenes covered in the experiments.
@@ 5.4 Ablation Study
We study different strategies of environment modeling for NVS in-the-wild, namely leaving the blank pixels with zeros, filling the blanks with a default learnable feature (as is done in RPBG), and wrapping the points with a sphere of $10^{6}$ points (equivalent to an environment map in [39]). The quantitative results are in Tab. 3, where the strategy incorporated by RPBG boosts the performance by a large margin, with almost no additional overhead.
@@
Table 3: Ablation on environment modeling on two unbounded scenes, namely Caterpillar of T&T dataset [24] and sky of Free dataset [54]. #Points (×$\times 10^{6}$) represent the equivalent overheads when applying different strategies.
@@
Zeros 7.95 19.89 0.456 0.403 17.49 21.77 0.598 0.315
Learnable 7.95 21.78 0.687 0.289 17.49 24.81 0.866 0.199
Sphere 8.95 21.68 0.686 0.283 18.49 25.15 0.869 0.179
@@ 6 Discussion
We would like to shed some light on the effectiveness of RPBG. The first factor should be the point representation. The triangulated 3D points have already contained all the verified co-visibility that is required for NVS. Compared to RF-based methods, which often struggle at local minima, such triangulation-as-parameterization paradigm greatly lower the complexity of searching, and is unified across different scene types. Besides, the patch-wise rendering scheme mentioned in Sec. 4.4 also improves the perceptual quality of re-rendering while RF-based methods apply pixel/ray-wise volume rendering, where the correlation between pixels are not modeled explicitly. For more discussions on the insights, we strongly refer the readers to the Supplementary Material.
@@
An obvious problem with point-based re-rendering (RPBG and NPBG [1]) is that compared to lightweight RF-based variants [7, 32], they take more space to store the CNN parameters and the neural texture. Although it makes the scaling-up easier, the number of parameters grows linearly with the point cloud scale. Besides, due to the non-local neural renderer we employ, each point is encoded with visual context of a larger range, hindering the editability of RPBG. We have also noticed that, the CNN-based rendering scheme can lead to unsatisfactory temporal consistency especially when the triangulation is patchy so that we mainly count on the inpainting ability of the neural render for yielding re-renderings. Such flicker issue, as well as its potential solutions, has been discussed in [12].
@@ 7 Conclusion
In this paper, we present RPBG as a robust and practical alternative to NPBG, a baseline of point-based NVS methods, performing neural re-rendering on triangulated points. We analyze the key problems in NPBG, when attempting to generalize to more generic scenes other than the well-captured scans, and reform the pipeline to reveal the real potential of point-based graphics. Respectively motivated and inspired by RF-based methods and low-level image restoration methods, we reform the pipeline according to our analysis. By extensive experiments on diverse datasets, RPBG achieves stably superior results over state-of-the-art RF-based and point-based NVS methods, especially on the metrics with more attention paid to the nuances of human visual perception, without case-by-case parameterization across all scenes, indicating its robustness and generalizability.
Acknowledgments
This research was supported in part by JSPS KAKENHI Grant Numbers 24K22318, 22H00529, 20H05951, and JST-Mirai Program JPMJMI23G1.
@@ Appendix 0.A Further Discussions
As the framed area in Fig. 9(a) shows, the rendering noise is mainly caused by under-representation of RF, which is further due to the sparsity of input views (lack of ray intersections). RF-based methods aim to represent the target scene loyally, where each inquiry is supposed to be a frank reflection of local optical properties. In this way, RF-based methods render an image in pixels without considering the context information, establishing better pixel-to-pixel correspondence (thus higher PSNR).
@@
Figure 9: Comparison on Museum of T&T dataset [24] to showcase the typically different sources of noise due to the fundamental differences between different types of methods. Zoom in for best view.
@@ Appendix 0.B Dense Triangulation
Based on the sparse triangulation, we follow the view selection strategy in [60], and choose 4 neighboring images with the best co-visibility for each image. Then we estimate a depth map for each image, by aid of the top-4 neighboring images, with AA-RMVSNet [56]. The per-view depth maps are filtered and fused to obtain the final 3D point cloud. We select AA-RMVSNet for its high memory-efficiency that allows a large batch size and RPBG is supposed to work fine with other off-the-self MVS methods.
@@ Appendix 0.C Point Cloud
The detailed augmentation steps are as Algorithm 1.
Algorithm 1 Point Cloud Augmentation
1:Input: The point cloud $\{X\}$ to be augmented
2:for a given number of times do
3:     Sample one existing point $X=(x,y,z)$ randomly from $\{X\}$
4:     Form a 3D Gaussian distribution $G$ with its mean value $\mu=X$
5:     Sample a new point $X^{\prime}$ from $G$
6:end for
7:Train RPBG with the point cloud $\{X\}\cup\{X^{\prime}\}$
8:for each $X_{i}$ in $\{X^{\prime}\}$ do
9:     Retrieve its neural texture $\mathbf{T}(X_{i})$
10:     Approximate its pseudo density $\sigma_{i}=\sum|\mathbf{T}(X_{i})|$
11:     if $\sigma_{i}<\sigma_{\textrm{threshold}}$ then
12:         Discard $X_{i}$ from $\{X^{\prime}\}$
13:     end if
14:end for
15:Output: The augmented point cloud $\{X\}\cup\{X^{\prime}\}$
On the scene of Church, we study the impact of such strategy applied on sparsely triangulated points multiple times. As is shown in Tab. 4, the first round of sampling and pruning brings the largest performance gain, and a larger gain is observed when applying to the SfM-initialized triangulation, which is much sparser compared to the MVS-initialized one. Note that the augmentation is optional and thus not performed on well triangulated scenes for the sake of time only.
@@
Since we have demonstrated in the main paper that RPBG is able to perform re-rendering even with a randomly initialized point cloud taken as input. Empirically, we find that by applying the spatial pruning strategy to the random point cloud (by thresholding point-wise $\sigma>180$ in this case), the point cloud shrinks to a shape similar to the actual geometry, as is illustrated in Fig. 10(b). It suggests that when neurally re-rendering, the network is able to implicitly verify the occupancy of each rasterized point and if a point is observed with poor multi-view consistency, it is more likely to be considered as an invalid point. The attempt of pruning random points is considered as an extreme case explaining how the point cloud augmentation strategy of RPBG manages to alleviate the problem of patchy or erroneous triangulation.
@@
The RF-based methods are sensitive to dynamic objects and require either data pre-processing, e.g., masking by manual labeling and semantic segmentation, or modeling of such ambiguity or uncertainty [40], to aid the RF’s optimization. As for RPBG, the robustness against transient objects is trivially achieved since they are typically not reconstructed in SfM or MVS for not satisfying the static scene assumption. By experiments, we discover that RPBG is robust to such dynamic objects and able to automatically such objects in the training views when re-rendering (Fig. 11), which suggests that multi-view consistency is implicitly enforced during training and the renderer tends to restore the most consensual re-rendering.
@@
As RPBG is a point-based pipeline, where an explicit 3D geometry is adopted for re-rendering, similar to previous point-based alternatives [1, 68], it allows certain scene editing and manipulation. We give two examples in Fig. 12. By removing the points, along with the point-bounded features, RPBG manages to re-render the edited scene, yet with some artifacts observed. It is because in RPBG, we enhance the context exchange among rasterized points by DAC, where each point does not solely represent its local optical property. Besides, it is also observed that FFC [17] may lead to repetitive artifacts at incomplete regions, as also can be found in the inpainted images by LaMa [44].
@@ Appendix 0.D More Quantitative Results
ScanNet++ Benchmark
We also evaluate RPBG on the public benchmark of ScanNet++ [61] (Novel View Synthesis on DSLR Images). ScanNet++ contains a wide variety of indoor scenes that are challenging for novel view synthesis for glossy and reflective materials and unseen poses captured independently of the training trajectory. The results are shown in Tab. 7. Note that the scores are all retrieved from the leaderboard. RPBG outperforms all the baselines listed by the benchmark, with a particular good perceptual quality (LPIPS).